find the maximum value of z=6x-3x^2+2y
(subject to y-x^2=2
if condition changes to y-x^2=3, how much will z increase?
Answer #1

The maximization problem is:
\small max \ 6x-3x^2+2y \ s.t. \ y-x^2=2
Setting up the Lagrange:
\small \mathcal{L}=6x-3x^2+2y+\lambda(y-x^2-2)
The FOCs are:
\small \frac{\partial \mathcal{L}}{\partial x}=6-6x-2\lambda x=0
\small \frac{\partial \mathcal{L}}{\partial y}=2+\lambda=0
\small y-x^2=2
Dividing the first two FOCs:
\small \frac{6-6x}{2}=\frac{2\lambda x}{-\lambda}\rightarrow x=3
Substituting this into the third FOC:
\small y-3^2=2\rightarrow y=11
The solution to the maximization problem is:
\small (x,y)=(3,11)

If the condition changes to \small y-x^2=3 ,
All the FOCs will remain the same except the last one. The last FOC will change to:
\small y-x^2=3
Since the first two FOCs do not change, the solution to the x coordinate will not change. Substituting the value of x into the new FOC:
\small y-3^2=3\rightarrow y=12
The solution to the new maximization problem is:
\small (x,y)=(3,12)